Using identities, evaluate. 
102^2 ???????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

102^2= (100 + 2)^2
= 100^2+ 400 + 2^2
= 10000 + 400 + 4
= 10404

Thanks